Now discuss this quadratic equation and determine the number and nature of the solutions, using the determine whether the solutions are real or complex?
3x2 - 2x - 1 = 0

Answer by AAfter.Com(18)

3x2 - 2x - 1 = 0
Diterminant D = (-2)^2 - 4(3)(-1)
= 4 + 12 = 16 > 0
Hence, the roots of the given equation are positive and real numbers.